OpenStudy (anonymous):

4 question math help! 1) y=6-x and y=x-2 a (2,4) b (-4,2) c (4,2) d no solution 2) y=2x-1 and y=x+3 a (4,7) b (7,4) c ( -7,-4) d infinite solutions 3) y=4x and y+x=5 a (-4,1) b (1,4) c ( -3,2) d (2,3) 4) What will the graph look like for a system of equations that has no solution? a) The lines will be perpendicular. b) The lines will cross at one point. c) Both equations will form the same line d) the lines will be parallel.

OpenStudy (mathmale):

What methods could you use to solve #1? 1) y=6-x and y=x-2 a (2,4) b (-4,2) c (4,2) d no solution Hints: among the many available methods, you could use 1) elimination by substitution, 2) elimination by addition and subtraction, or 3) graphing.

OpenStudy (anonymous):

I have no idea how to do these... I have been working on these 4 questions for the past 4 hours.. I have had multiple people tell me how to do them.. I will just never understand them.

OpenStudy (mathmale):

I'm sorry...you must be really frustrated. But please think back to what you've seen and heard in class, or what you've read in your online learning materials. What do you recall about "solving systems of linear equations?"

OpenStudy (mathmale):

I am going to demo how to solve this particular system in 2 different ways, hoping that this will ring a bell with you (help you to remember material that you already know): If the system is y=6-x and y=x-2 then we can set y=6-x equal to y=x-2; we thereby eliminate y and get 6-x=x-2. If we add x to both sides, we get 6=2x-2. If we now add 2 to both sides, we get 8=2x and solving that for x results in x=4. Can you now find y?

OpenStudy (anonymous):

okay 1 sec I'm going to write this down...

OpenStudy (mathmale):

Your system is y=6-x and y=x-2. You could insert x=4 into either or both of these equations to check your result.

OpenStudy (anonymous):

what do you mean by insert it?

OpenStudy (mathmale):

We got x=4 earlier. We want to find y. So, taking the equation y=6-x and substituting 4 for x, we get y=6-4=2. So, if x = 4, y=2. To be doubly sure that our result is correct, we could substitute x = 4 into the second equation also: y=x-2. substituting x=4, y=4-2 = 2. Same answer as before. Therefore, (4,2) IS the solution of the system of equations y=6-x and y=x-2 .

OpenStudy (anonymous):

Ohh okayy... I think I am actually starting to get it now! THank you so much! ^_^

OpenStudy (mathmale):

Never give up on yourself!!!! Glad that this is starting to make more sense. I mentioned 3 ways of solving this system of linear equations: y=6-x and y=x-2 1. elimination by substitution 2. elimination by adding and subtracting 3. graphing It would be well worth learning to do all three of these. You may even want to learn other methods later: determinants, matrices and the like.

OpenStudy (mathmale):

Excuse me: I need to get off the 'Net now. Good luck to you!